# Interfacial spin-orbit coupling driven enhancement of superconductivity by parallel magnetic field

### v1

@@ Abstract
Figure 4: Sketch of a hollow superconducting cylinder placed into external magnetic field ${\bf H_{0}}$. Broken inversion symmetry at the internal surface results in emergence of Rashba spin-orbit interaction on this interface.
The typical dependencies $\Delta T_{c}(H_{0})$ for different SOC parameters are shown in Fig.3. Note that for $L\sim 0.2\xi$ the parameter $p_{soc}=12$ correspond to SOC constant $s_{soc}\sim 0.12$, which is reasonable value. We also assume strong spin-orbit scattering in the system, which decrease paramagnetic contribution to $\Delta T_{c}$.
Our calculations show, that for reasonable choice of parameters the superconducting critical temperature $T_{c}$ of a thin ($L\ll\xi$) superconducting film with interfacial Rashba SOC can be enhanced by external magnetic field in wide range of the fields, see. Fig.3. This enhancement is caused by Lifshitz invariant in the free energy functional Eq.1. The results (see. Fig.3) are in qualitative agreement with the experimental data obtained in Refs.[Jeffrey Gardner et al., 2011; Asaba et al., 2018]. Note that we expect, that in all systems experimentally studied in Refs.[Jeffrey Gardner et al., 2011; Asaba et al., 2018] spin-orbit scattering is strong.
Interfacial Rashba SOC can also modify Little-Parks oscillations of $T_{c}$ [see Fig.5]. To show this, let us consider a hollow superconducting cylinder having internal (external) radius $R_{1}$ ($R_{2}$) with broken inversion symmetry at the internal surface, resulting in Rashba spin-orbit interaction, placed into external magnetic field ${\bf H}_{0}=H_{0}{\bf e}_{z}$, see Fig.4. We assume the cylinder to be thin-walled, i.e. $\Delta R=(R_{2}-R_{1})\ll\xi(0)$. In such the case the order parameter in cylindrical coordinate system $(r,\varphi,z)$ read as $\psi({\bf r})=\psi_{0}\exp{(il\varphi)}$, where $l$ is integer angular momentum. As a result, the relative change of the critical temperature $\tau=(T_{c}-T_{c0})/T_{c0}$ is the following:
@@
Here $\mu=\Phi/\Phi_{0}$, where $\Phi=\pi H_{0}R_{1}^{2}$ is external magnetic flux through the cylinder, $\rho=R_{1}/\xi(0)$, $\delta=\Delta R/R_{1}$ and $\tilde{A}=\sqrt{b}\xi/R_{1}$.
Figure 5: The superconducting critical temperature $T_{c}$ of a hollow thin-walled cylinder as a function of external magnetic flux $\Phi$. Here Rashba SOC parameter $s_{soc}=0.12$ for red curve, while SOC is absent for blue and green ones. The value of paramagnetic effect (PME) is $\tilde{\chi}=0.24$ for red and green curve. The blue curve demonstrates Little-Parks (LP) oscillations without both PME and SOC. For all curves $R_{1}/\xi(0)=4$, $\Delta R/\xi(0)=0.2$ and $b\sim 1$.
The first term in Eq.(13) describes usual Little-Parks oscillations Parks and Little (1964). The second one is responsible for suppression of $T_{c}$ due to paramagnetic effect. In contrary, the last term in Eq.(13), caused by the combination of interfacial Rashba SOC and external magnetic field, lead to the enhancement of $T_{c}$. As a result, the envelope function of $T_{c}(\Phi)$ dependence demonstrates nonmonotonic behaviour: it grows initially, reaches the maximum and then declines. This nonmonotonic envelope is superimposed by Little-Parks like oscillations [see Fig.5]. As a result, for reasonable choice of parameters the $T_{c}$ of hollow thin-walled cylinder could be enhanced, but not suppressed, by external magnetic flux.
In conclusion, we show that the critical temperature of thin superconducting film with broken inversion symmetry and strong SOC in the bulk can be enhanced by external magnetic field, see Fig.3. This effect is caused by Lifshitz invariant in GL functional, which results from combination of external magnetic field and interfacial Rashba SOC, appearing, on its tern, due to lack of structural inversion symmetry combined with bulk SOC. The results are in qualitative agreement with the experimental data Jeffrey Gardner et al. (2011); Asaba et al. (2018). The same mechanism can cause the modification of Little-Parks oscillations in cylindrical geometry, see Fig.5.
This work is supported by Russian Science Foundation, (grant #22-72-00083) in part related to the calculation of the critical field of thin film and Ministry of Science and Higher Education of the Russian Federation within the framework of state funding for the creation and development of World-Class Research Center (WCRC) “Digital biodesign and personalized healthcare”, grant no. 075-15-2022-304, in part related to the calculation of the influence of the spin-orbit interaction on the Little-Parks oscillations. Zh.D. acknowledges the support from ”Basis” foundation.

### v2

@@ Abstract
The typical dependencies $\Delta T_{c}(H_{0})$ for different SOC parameters are shown in Fig.3. Note that for $L\sim 0.2\xi$ the parameter $p_{soc}=12$ correspond to SOC constant $s_{soc}\sim 0.12$, which is reasonable value. We also assume strong spin-orbit scattering in the system, which decrease the paramagnetic contribution to $\Delta T_{c}$.
Our calculations show, that for reasonable choice of parameters the superconducting critical temperature $T_{c}$ of a thin ($L\ll\xi$) superconducting film with interfacial Rashba SOC can be enhanced by external magnetic field in wide range of the fields (see. Fig.3). This enhancement is caused by Lifshitz invariant in the GL functional Eq.(1). The results (see. Fig.3) are in qualitative agreement with the experimental data obtained in Refs.[Jeffrey Gardner et al., 2011; Asaba et al., 2018]. Note that we expect spin-orbit scattering to be strong in all systems experimentally studied in Refs.[Jeffrey Gardner et al., 2011; Asaba et al., 2018].
Figure 4: Sketch of a hollow superconducting cylinder placed into magnetic field. The broken inversion symmetry at the internal surface results in the emergence of Rashba SOC.
Interfacial Rashba SOC can also modify Little-Parks oscillations of $T_{c}$ (see Fig.5). To show this, let us consider a hollow superconducting cylinder having internal (external) radius $R_{1}$ ($R_{2}$) with broken inversion symmetry at the internal surface, resulting in Rashba SOC, placed into the external magnetic field ${\bf H}_{0}=H_{0}{\bf e}_{z}$ (see Fig.4). We assume the cylinder to be thin-walled, i.e. $\Delta R=(R_{2}-R_{1})\ll\xi(0)$. In such the case the order parameter in cylindrical coordinate system $(r,\varphi,z)$ reads as $\psi({\bf r})=\psi_{0}\exp{(il\varphi)}$, where $l$ is integer angular momentum. As a result, the relative change of the critical temperature $\tau=(T_{c}-T_{c0})/T_{c0}$ is the following:
@@
Here $\mu=\Phi/\Phi_{0}$, where $\Phi=\pi H_{0}R_{1}^{2}$ is the external magnetic flux through the cylinder, $\rho=R_{1}/\xi(0)$, $\delta=\Delta R/R_{1}$ and $\tilde{A}=\sqrt{b}\xi/R_{1}$.
The first term in Eq.(13) describes usual Little-Parks oscillations Parks and Little (1964). The second one is responsible for the suppression of $T_{c}$ due to the paramagnetic effect. In contrary, the last term in Eq.(13), caused by the combination of interfacial Rashba SOC and external magnetic field, leads to the enhancement of $T_{c}$. As a result, the envelope function of the $T_{c}(\Phi)$ dependence demonstrates nonmonotonic behaviour: it grows initially, reaches the maximum and then declines. This nonmonotonic envelope is superimposed by Little-Parks like oscillations (see Fig.5). As a result, for reasonable choice of parameters the $T_{c}$ of a hollow thin-walled cylinder could be enhanced, but not suppressed, by external magnetic flux.
Figure 5: The superconducting critical temperature $T_{c}$ of a thin-walled cylinder as a function of the external magnetic flux $\Phi$. Here Rashba SOC parameter $s_{soc}=0.1$ for the red curve, while SOC is absent for the blue and green ones. The paramagnetic parameter is $\alpha_{p}=0.06$ for the red and green curves. The blue curve demonstrates Little-Parks (LP) oscillations without both paramagnetic effect (PME) and SOC. For all curves $R_{1}/\xi(0)=4$, $\Delta R/\xi(0)=0.2$ and $b\sim 1$.
In conclusion, we show that the critical temperature of a thin superconducting film with broken inversion symmetry and strong SOC in the bulk can be enhanced by external magnetic field (see Fig.3). This effect is caused by Lifshitz invariant in GL functional, which results from combination of external magnetic field and interfacial Rashba SOC, appearing, on its tern, due to the lack of structural inversion symmetry combined with bulk SOC. The results are in qualitative agreement with the experimental data Jeffrey Gardner et al. (2011); Asaba et al. (2018). The same mechanism can cause the modification of Little-Parks oscillations in cylindrical geometry (see Fig.5).
The authors express their gratitude to V.G. Kogan for his valuable correspondence. This work is supported by Russian Science Foundation (grant #22-72-00083) in part related to the calculation of the critical field of thin film and Ministry of Science and Higher Education of the Russian Federation within the framework of state funding for the creation and development of World-Class Research Center (WCRC) “Digital biodesign and personalized healthcare” (grant #075-15-2022-304) in part related to the calculation of the influence of the spin-orbit interaction on the Little-Parks oscillations. Zh.D. acknowledges the support from ”Basis” foundation.